Find the hydrostatic and absolute pressure exerted on an object that is 4.45m below the surface of gasoline. The density of gasoline is 680 kg/m3
1
Expert's answer
2020-03-26T12:17:46-0400

hydrostatic pressure is determined as follows

"P.hydrostatic = \rho * g * h = 680 * 9.81 * 4.45 = 29,685 Pa"

Absolute pressure is the total of atmospheric and hydrostatic pressure

"P.absolute = P.hydrstatic + P.atmosphere = 29,685 + 100,000 = 129,685 Pa"